Mia Benson invested her summer earnings of R3 000 in a savings account which pays 2.5% interest. How much will this amount to in 6 months?

1
Expert's answer
2021-11-24T15:09:49-0500

"FV=PV(1+r) 
^n\\"

after 6 months:

"FV=3000(1+0.025) \n^{1\/2}\n =R 3037.27"